The interests earned on an investment varies directly with interest rate. If a 10% rate gives 200$, the interest rate that gives 300$ is..
Answer by checkley77(12844)

.10/200=x/300
200x=.1*300
200x=30
x=30/200
x=.15 or 15% interst rate yields the $300.